# Modelling Cosmic-Ray Transport: Magnetized versus Unmagnetized Motion in Astrophysical Magnetic Turbulence

### v2

@@ 2.2 Test Particle Setup
In order to investigate scattering and transport behaviour of particles in the magnetized and unmagnetized regimes, we consult the geometric picture of Lemoine (2023) and Kempski et al. (2023) based on the field-line curvature
@@
where $\kappa r_{g}\ll 1$ corresponds to the magnetized case, $\kappa r_{g}\gg 1$ corresponds to the unmagnetized case, and for $\kappa r_{g}\sim 1$ order-unity variations of the magnetic moment
@@
Further, we record spatial mean squared displacements (MSD) of particles
@@
conditional on the magnetization criteria $\bar{\kappa}\bar{r}_{g}<1$ and $\bar{\kappa}\bar{r}_{g}>1$, as well as an unconditional baseline. We assume stationarity $\langle\|\boldsymbol{X}_{\tau}-\boldsymbol{X}_{0}\|^{2}\rangle=\langle\|\boldsymbol{X}_{t+\tau}-\boldsymbol{X}_{t}\|^{2}\rangle$, expressed by the local time scale $\tau$. The MSD, obtained from the displacement of individual particles, measures the time-dependent spread of the underlying particle distribution function, and classifies the motion as super-, normal or sub-diffusive, depending on $\langle\Delta X^{2}_{\tau}\rangle\sim\tau^{\alpha}$ scaling with $\alpha>1$=1=1 or <1<1. For $\alpha=1$, the classical diffusion coefficient is given by $D_{\infty}=\lim_{\tau\to\infty}\langle\Delta X^{2}_{\tau}\rangle/2\tau$ (Metzler & Klafter, 2000).
The unconditional MSD should be treated with care, because it averages over many different length scales, thus hiding relevant physical processes, which we address by conditioning on the magnetization criterion. Further, since our simulation box only contains one flow correlation cell, and the parallel scale of the magnetic field $k_{\parallel}$ is comparable to the box size, $D_{\infty}$ converges only after particles traverse the periodic domain multiple times. In doing so, particles likely re-enter the box at a different position, and thus sample different regions of the magnetic field, which results in a meaningful albeit biased value for $D_{\infty}$.
In order to obtain the conditional statistics, we simulated for each $\hat{\omega}_{g}$, in each of the 8 statistically independent MHD snapshots, 400 000400\,000 independent test-particle trajectories for $1000$ gyro periods. The unconditional baseline MSD is, per snapshot, based on 40 00040\,000 independent test-particle trajectories with lengths of 10 00010\,000 gyro periods. This ensures that the tails of the joint density $p(\bar{\kappa},\bar{r}_{g})$ are well resolved. However, the conditional MSD at large time scales is hard to resolve properly, even with increased sample sizes. This matter is discussed in more detail in Section 3.2.
@@ 2.3 MHD Simulation Results
Figure 4: Magnetic field lines coloured by field strength $B$ and test-particle trajectories coloured by pitch angle cosine $\mu$ in the flux tube (top) and one of the plasmoids (bottom) from figure 3. In the coherent flux tube, particles are closely bound to field lines with occasional large-scale mirroring. The plasmoid exhibits highly tangled field lines and effectively confines particles with a mixture of small-scale mirroring and unmagnetized scattering. The magnetic correlation scale is indicated for reference.
We start by inspecting the magnetic field snapshots extracted from the statistically saturated phase of the MHD simulation. Figure 2 shows isosurfaces of the magnetic field strength $B$ and the current density magnitude $j=\|\nabla\times\boldsymbol{B}\|$ at different zoom levels in the simulation box. The isosurfaces of $B$ reveal characteristic structures of the saturated fluctuating dynamo (Miller, 2019; Seta et al., 2020), which can be described as long flattened tubes with length $l$, width $w$, and height $h$, ordered as $l\gg w>h$. As indicated in figure 4, they contain coherent bundles of field lines and approximately represent flux surfaces. Thus, henceforth, we refer to these coherent structures as flux tubes. The connection between field strength $B$ and coherent geometry with small curvatures $\kappa$ is formally expressed by the anti-correlation between the two quantities $B\sim\kappa^{-1/2}$ (Schekochihin et al., 2004).
At the outermost zoom level in figure 2, we recognize that most of the magnetic energy is distributed intermittently throughout the domain, concentrated in a few intense flux tubes with almost circular cross-sections, which extend up to the flow correlation scale $l\lesssim L_{u}$. These structures are sometimes observed wrapped up by intense current sheets. Further zooming into apparently quiet regions reveals a population of smaller, less intense, flattened flux tubes, organized in a tightly folded pattern (Schekochihin et al., 2004) with embedded current sheets. The networks of flux tubes and current sheets are dual to each other. Outside of coherent flux tubes, e.g. in the debris of disrupted structures, field lines tend to be chaotically tangled and highly disorganized. Additionally, reconnection (either due to resistive diffusion or due to tearing) of folded flux tubes is observed to seed small-scale plasmoids (Galishnikova et al., 2022). This diverse picture is illustrated by a slice plot through an intense flux tube and its surroundings in figure 3.
The conventionally computed correlation wavenumber $k_{B,\mathrm{corr}}$ does not adequately reflect the geometric structure of the magnetic field, as indicated by its rather large value $O(10\,k_{\mathrm{forcing}})$, because the intermittent and highly correlated flux tubes are averaged out. It is thus instructive to also report the characteristic geometric scales $k_{\parallel}$$k_{\boldsymbol{B}\times\boldsymbol{j}}$ and $k_{\boldsymbol{B}\cdot\boldsymbol{j}}$, which are sensitive to the local direction of the magnetic field (see table 2; Schekochihin et al., 2004; Galishnikova et al., 2022). For $h=2$, these values confirm elongated tube-like structures with length $l_{\parallel}\approx L_{u}$ and an approximate aspect ratio of $1:6:6$. On the other hand, the flow $\boldsymbol{u}$ exhibits a lower degree of intermittency compared to $\boldsymbol{B}$, so the correlation wavenumber is more suitable to characterise its correlation structure. The small discrepancy between $k_{u,\mathrm{corr}}$ and $k_{\mathrm{forcing}}$ results from the presence of a broadband turbulent energy spectrum (Seta et al., 2020).
@@ 2.4 Test Particle Results
Figure 4 shows examples of test-particle trajectories. Particles are clearly magnetized with $\bar{\kappa}\bar{r}_{g}<1$ inside the intense flux tube, where they closely follow magnetic field lines. Due to large-scale variations of the illustrated flux tube, particles undergo occasional mirroring events, which appear as sudden reversals of direction. These rare reversals imply a long parallel MFP. Outside of the flux tube, particle motion is unmagnetized with $\bar{\kappa}\bar{r}_{g}>1$ and appears much more erratic, as the particles bounce chaotically and with a short MFP through the tangled field. An additional interesting case is illustrated by the small-scale plasmoid, which appears to be an efficient device for trapping particles. However, since they appear to have a sub-dominant contribution in our statistics, and due to numerical concerns addressed in Section 4.2, we focus here on the modelling of magnetized and unmagnetized motion and leave a detailed study on the effect of plasmoids on particle transport for later work.
@@
Figure 5: (a) Average of the relative magnetic moment variation $\overline{\!{\delta M}}/\,\overline{\!{M}}$ conditional on particle gyro radius $\bar{r}_{g}$ and field-line curvature $\bar{\kappa}$. All recorded quantities are gyro-averaged. The colour scale is centred at $\overline{\!{\delta M}}/\,\overline{\!{M}}=1$, where particles can be considered magnetized for smaller variations and unmagnetized for larger variations. Further, the colour scale is capped to $\log_{10}\,\overline{\!{\delta M}}/\,\overline{\!{M}}\in(-0.6,0.6)$ in order to highlight the transition region. This transition region is compared with the magnetization criterion $\bar{\kappa}\bar{r}_{g}\sim 1$ expected from the field-line curvature picture. The joint density $p(\bar{\kappa},\bar{r}_{g})$ is indicated for reference. (b) The conditional average $\langle\,\overline{\!{\delta M}}/\,\overline{\!{M}}|\bar{\kappa}\bar{r}_{g}\rangle$ also shows the transition from predominantly magnetized and to predominantly unmagnetized motion as $\bar{\kappa}\bar{r}_{g}$ increases, although the joint density $p(\,\overline{\!{\delta M}}/\,\overline{\!{M}},\bar{\kappa}\bar{r}_{g})$ reveals some uncertainty of this criterion.
@@
Figure 6: (a) MSD of the test particles, depending on the time lag $\tau$, revealing initial ballistic propagation, transient subdiffusion and asymptotic diffusive behaviour. Notably, diffusion only occurs for mean square displacements beyond the simulation box size. (b) Conditional MSD for magnetized and unmagnetized transport for a selected test-particle energy, as well as the respective relative bin counts. Longer consecutively magnetized or unmagnetized segments have a lower probability than shorter ones, which leads to a systematically smaller sample size for the conditional averages at larger time scales $\tau$. We describe magnetized transport by compound subdiffusion, and unmagnetized transport by a Langevin equation. The discrepancy between magnetized data and model are likely due to this bias at large $\tau$, which is addressed when tuning the combined model to the unconditional MSD, resulting in good agreement.
@@
These observations are substantiated by the conditional statistics given by equations (8) and (9). Firstly, figure 5 confirms that particles are magnetized in the sense that their magnetic moments are weakly conserved, $\overline{\!{\delta M}}/\,\overline{\!{M}}<1$, when their gyro radii are smaller than the experienced field-line curvature radius, i.e. $\bar{\kappa}\bar{r}_{g}<1$ (Lemoine, 2023). If this condition is violated, i.e. $\bar{\kappa}\bar{r}_{g}>1$, large variations of the magnetic moment are expected. Although $\bar{\kappa}\bar{r}_{g}=1$ does not constitute an exact decision boundary, we consider it as a useful heuristic, because our ensuing results remain qualitatively robust under minor refinements of this condition. Appendix B shows results for an alternative magnetization criterion based on the perpendicular reversal scale $\kappa_{\perp}$ introduced by Kempski et al. (2023), which can be approximated as $\bar{\kappa}_{\perp}\bar{r}_{g}^{1/2}\sim 30$.
@@
Secondly, figure 6 shows the unconditional MSD for our considered particles, as well as the conditional observations for $\hat{\omega}_{g}=181.019$. The unconditional data reveals three time scales of transport: particles move ballistically on short time scales, followed by a transient subdiffusive phase, before becoming asymptotically diffusive. As noted in Section 2.2, convergence to normal diffusion occurs only after the MSD reaches the box size, which means that particles may experience the same structures repeatedly due to the periodic boundary conditions. We address this issue partly by modelling the transport behaviour on short and intermediate time scales, where the bias of the periodic domain is negligible.
@@
The conditional data supports two distinct transport behaviours, where magnetized particles exhibit much longer MFPs compared to unmagnetized particles. Also indicated are the relative bin counts contributing to the conditional averages. This number strongly decreases for both cases for larger time scales, because the probability of finding an uninterrupted conditional trajectory segment decreases with its desired length. The conditional averages at large time scales are thus likely dominated by a few very intense flux tubes, implying a systematic bias caused by our heuristic magnetization criterion $\bar{\kappa}\bar{r}_{g}\sim 1$. Our methodology for modelling the conditional and unconditional data, as well as addressing this bias, is presented in Section 3.
@@ 3 Stochastic Transport Model
Motivated by our previous findings, we present a simplified stochastic model for the transport of charged test particles, resulting from a competition between magnetized and unmagnetized behaviour. We assume that magnetized behaviour is dictated by the most intense flux tubes which extend up to the correlation scale of the turbulence $\lambda_{\mathrm{fl}}\lesssim L_{u}$. Particles are strictly bound to those field lines, but may undergo pitch-angle scattering with rate $\theta_{\mu}$. This scattering represents large-scale mirror interactions on the scale of the flux tubes, and thus, its MFP $\lambda_{\mu}\sim c/\theta_{\mu}$ is expected to be long and comparable to the flux tube scale, i.e., $\lambda_{\mu}\sim\lambda_{\mathrm{fl}}$. Particles remain in the magnetized transport state for some mean duration $t_{\mu}^{\ast}$, which represents the mean time until a scattering event due to sharp field-line curvature occurs.
On the other hand, unmagnetized behaviour results from an interplay between the intrinsic particle momentum and random scattering due to highly tangled field lines, where no well-defined mean field is seen by the particle, so we expect a relatively short MFP and high scattering rate $\theta_{\mathrm{scatter}}\sim c/\lambda_{\mathrm{scatter}}$. Particles wander around the domain for some mean duration $t^{\ast}_{\mathrm{scatter}}$ until they encounter a strong flux tube and become magnetized again.
@@
The mean durations parametrize escape-time probability distributions $p_{\mu}(t|t^{\ast}_{\mu})$ and $p_{\mathrm{scatter}}(t|t^{\ast}_{\mathrm{scatter}})$ for both transport modes. They may be related to the distribution of field-line curvature, and from the dominance of low curvatures we expect the ordering
@@
We assume here that each kind of motion (specifically field-line wandering, magnetized mirroring and unmagnetized scattering) is sufficiently characterized by its respective MFP $\lambda$ and thus adequately modelled by a Langevin equation, where both initial ballistic and asymptotic diffusive behaviour are resolved. The emergent combined transport behaviour can then be studied as a competition between the involved MFPs and mean durations. Further, the combination of two simple Langevin-like descriptions readily leads to compound subdiffusion, and thus a plausible explanation for the transient subdiffusion shown by the test particles. We emphasize that qualitative descriptions are linked to the relative sizes of the various MFPs, for instance “coherent motion” or “large-scale mirroring” correspond to $\lambda_{\mu}\sim L_{u}$, while “chaotic motion” corresponds to $\lambda_{\mathrm{scatter}}\ll L_{u}$.

### v3

@@ 2.2 Test particle set-up
To investigate scattering and transport behaviour of particles in the magnetised and unmagnetised regimes, we consult the geometric picture of Lemoine (2023) and Kempski et al. (2023) based on the field-line curvature
@@
where $\kappa r_{g}\ll 1$ corresponds to the magnetised case, $\kappa r_{g}\gg 1$ corresponds to the unmagnetised case and for $\kappa r_{g}\sim 1$ order-unity variations of the magnetic moment,
@@
Further, we record spatial mean squared displacements (MSDs) of particles
@@
conditional on the magnetisation criteria $\bar{\kappa}\bar{r}_{g}<1$ and $\bar{\kappa}\bar{r}_{g}>1$, as well as an unconditional baseline. We assume stationarity $\langle\|\boldsymbol{X}_{\tau}-\boldsymbol{X}_{0}\|^{2}\rangle=\langle\|\boldsymbol{X}_{t+\tau}-\boldsymbol{X}_{t}\|^{2}\rangle$, expressed by the local time scale $\tau$. The MSD, obtained from the displacement of individual particles, measures the time-dependent spread of the underlying particle distribution function, and classifies the motion as super-, normal or sub-diffusive, depending on $\langle\Delta X^{2}_{\tau}\rangle\sim\tau^{\alpha}$ scaling with $\alpha>1$=1=1 or <1<1. For $\alpha=1$, the classical diffusion coefficient is given by $D_{\infty}=\lim_{\tau\to\infty}\langle\Delta X^{2}_{\tau}\rangle/2\tau$ (Metzler & Klafter, 2000).
The unconditional MSD should be treated with care, because it averages over many different length scales, thus hiding relevant physical processes, which we address by conditioning on the magnetisation criterion. Further, since our simulation box only contains one flow correlation cell, and the parallel scale of the magnetic field $k_{\parallel}$ is comparable to the box size, $D_{\infty}$ converges only after particles traverse the periodic domain multiple times. In doing so, particles likely re-enter the box at a different position and thus sample different regions of the magnetic field, which results in a meaningful albeit biased value for $D_{\infty}$.
To obtain the conditional statistics, we simulated for each $\hat{\omega}_{g}$, in each of the eight statistically independent MHD snapshots, 400 000400\,000 independent test-particle trajectories for $1000$ gyro periods. The unconditional baseline MSD is, per snapshot, based on 40 00040\,000 independent test-particle trajectories with lengths of 10 00010\,000 gyro periods. This ensures that the tails of the joint density $p(\bar{\kappa},\bar{r}_{g})$ are well resolved. However, the conditional MSD at large time scales is hard to resolve properly, even with increased sample sizes. This matter is discussed in more detail in Section 3.2.
@@ 2.3 MHD simulation results
We start by inspecting the magnetic field snapshots extracted from the statistically saturated phase of the MHD simulation. Figure 2 shows isosurfaces of the magnetic field strength $B$ and the current density magnitude $j=\|\nabla\times\boldsymbol{B}\|$ at different zoom levels in the simulation box. The isosurfaces of $B$ reveal characteristic structures of the saturated fluctuating dynamo (Miller, 2019; Seta et al., 2020), which can be described as long flattened tubes with length $l$, width $w$ and height $h$, ordered as $l\gg w>h$. As indicated in figure 4, they contain coherent bundles of field lines and approximately represent flux surfaces. Thus, henceforth, we refer to these coherent structures as flux tubes. The connection between field strength $B$ and coherent geometry with small curvatures $\kappa$ is formally expressed by the anti-correlation between the two quantities $B\sim\kappa^{-1/2}$ (Schekochihin et al., 2004).
At the outermost zoom level in figure 2, we recognise that most of the magnetic energy is distributed intermittently throughout the domain, concentrated in a few intense flux tubes with almost circular cross-sections, which extend up to the flow correlation scale $l\lesssim L_{u}$. These structures are sometimes observed wrapped up by intense current sheets. Further zooming into apparently quiet regions reveals a population of smaller, less intense, flattened flux tubes, organised in a tightly folded pattern (Schekochihin et al., 2004) with embedded current sheets. The networks of flux tubes and current sheets are dual to each other. Outside of coherent flux tubes, e.g. in the debris of disrupted structures, field lines tend to be chaotically tangled and highly disorganised. Additionally, reconnection (either due to resistive diffusion or due to tearing) of folded flux tubes is observed to seed small-scale plasmoids (Galishnikova et al., 2022). This diverse picture is illustrated by a slice plot through an intense flux tube and its surroundings in figure 3.
The conventionally computed correlation wavenumber $k_{B,\mathrm{corr}}$ does not adequately reflect the geometric structure of the magnetic field, as indicated by its rather large value $O(10\,k_{\mathrm{forcing}})$, because the intermittent and highly correlated flux tubes are averaged out. It is thus instructive to also report the characteristic geometric scales $k_{\parallel}$$k_{\boldsymbol{B}\times\boldsymbol{j}}$ and $k_{\boldsymbol{B}\cdot\boldsymbol{j}}$, which are sensitive to the local direction of the magnetic field (see table 2; Schekochihin et al., 2004; Galishnikova et al., 2022). For $h=2$, these values confirm elongated tube-like structures with length $l_{\parallel}\approx L_{u}$ and an approximate aspect ratio of $1:6:6$. However, the flow $\boldsymbol{u}$ exhibits a lower degree of intermittency compared with $\boldsymbol{B}$, so the correlation wavenumber is more suitable to characterise its correlation structure. The small discrepancy between $k_{u,\mathrm{corr}}$ and $k_{\mathrm{forcing}}$ results from the presence of a broadband turbulent energy spectrum (Seta et al., 2020).
@@
Figure 4: Magnetic field lines coloured by field strength $B$ and test-particle trajectories coloured by pitch angle cosine $\mu$ in the flux tube (top) and one of the plasmoids (bottom) from figure 3. In the coherent flux tube, particles are closely bound to field lines with occasional large-scale mirroring. The plasmoid exhibits highly tangled field lines and effectively confines particles with a mixture of small-scale mirroring and unmagnetised scattering. The magnetic correlation scale is indicated for reference.
@@ 2.4 Test particle results
Figure 4 shows examples of test-particle trajectories. Particles are clearly magnetised with $\bar{\kappa}\bar{r}_{g}<1$ inside the intense flux tube, where they closely follow magnetic field lines. Due to large-scale variations of the illustrated flux tube, particles undergo occasional mirroring events, which appear as sudden reversals of direction. These rare reversals imply a long parallel MFP. Outside of the flux tube, particle motion is unmagnetised with $\bar{\kappa}\bar{r}_{g}>1$ and appears much more erratic, as the particles bounce chaotically and with a short MFP through the tangled field. An additional interesting case is illustrated by the small-scale plasmoid, which appears to be an efficient device for trapping particles. However, since they appear to have a sub-dominant contribution in our statistics, and due to numerical concerns addressed in Section 4.2, we focus here on the modelling of magnetised and unmagnetised motion, and leave a detailed study on the effect of plasmoids on particle transport for later work.
@@
Figure 5: (a) Average of the relative magnetic moment variation $\overline{\!{\delta M}}/\,\overline{\!{M}}$ conditional on particle gyro radius $\bar{r}_{g}$ and field-line curvature $\bar{\kappa}$. All recorded quantities are gyro-averaged. The colour scale is centred at $\overline{\!{\delta M}}/\,\overline{\!{M}}=1$, where particles can be considered magnetised for smaller variations and unmagnetised for larger variations. Further, the colour scale is capped to $\log_{10}\,\overline{\!{\delta M}}/\,\overline{\!{M}}\in(-0.6,0.6)$ to highlight the transition region. This transition region is compared with the magnetisation criterion $\bar{\kappa}\bar{r}_{g}\sim 1$ expected from the field-line curvature picture. The joint density $p(\bar{\kappa},\bar{r}_{g})$ is indicated for reference. (b) The conditional average $\langle\,\overline{\!{\delta M}}/\,\overline{\!{M}}|\bar{\kappa}\bar{r}_{g}\rangle$ also shows the transition from predominantly magnetised and to predominantly unmagnetised motion as $\bar{\kappa}\bar{r}_{g}$ increases, although the joint density $p(\,\overline{\!{\delta M}}/\,\overline{\!{M}},\bar{\kappa}\bar{r}_{g})$ reveals some uncertainty of this criterion.
@@
Figure 6: (a) MSD of the test particles, depending on the time lag $\tau$, revealing initial ballistic propagation, transient subdiffusion and asymptotic diffusive behaviour. Notably, diffusion only occurs for mean square displacements beyond the simulation box size. (b) Conditional MSD for magnetised and unmagnetised transport for a selected test-particle energy, as well as the respective relative bin counts. Longer consecutively magnetised or unmagnetised segments have a lower probability than shorter ones, which leads to a systematically smaller sample size for the conditional averages at larger time scales $\tau$. We describe magnetised transport by compound subdiffusion and unmagnetised transport by a Langevin equation. The discrepancy between magnetised data and model are likely due to this bias at large $\tau$, which is addressed when tuning the combined model to the unconditional MSD, resulting in good agreement.
@@
These observations are substantiated by the conditional statistics given by equations (8) and (9). First, figure 5 confirms that particles are magnetised in the sense that their magnetic moments are weakly conserved, $\overline{\!{\delta M}}/\,\overline{\!{M}}<1$, when their gyro radii are smaller than the experienced field-line curvature radius, i.e. $\bar{\kappa}\bar{r}_{g}<1$ (Lemoine, 2023). If this condition is violated, i.e. $\bar{\kappa}\bar{r}_{g}>1$, large variations of the magnetic moment are expected. Although $\bar{\kappa}\bar{r}_{g}=1$ does not constitute an exact decision boundary, we consider it as a useful heuristic, because our ensuing results remain qualitatively robust under minor refinements of this condition. Appendix B shows results for an alternative magnetisation criterion based on the perpendicular reversal scale $\kappa_{\perp}$ introduced by Kempski et al. (2023), which can be approximated as $\bar{\kappa}_{\perp}\bar{r}_{g}^{1/2}\sim 30$.
@@
Second, figure 6 shows the unconditional MSD for our considered particles, as well as the conditional observations for $\hat{\omega}_{g}=181.019$. The unconditional data reveals three time scales of transport: particles move ballistically on short time scales, followed by a transient subdiffusive phase, before becoming asymptotically diffusive. As noted in Section 2.2, convergence to normal diffusion occurs only after the MSD reaches the box size, which means that particles may experience the same structures repeatedly due to the periodic boundary conditions. We address this issue partly by modelling the transport behaviour on short and intermediate time scales, where the bias of the periodic domain is negligible.
@@
The conditional data supports two distinct transport behaviours, where magnetised particles exhibit much longer MFPs compared to unmagnetised particles. Also indicated are the relative bin counts contributing to the conditional averages. This number strongly decreases for both cases for larger time scales, because the probability of finding an uninterrupted conditional trajectory segment decreases with its desired length. The conditional averages at large time scales are thus likely dominated by a few very intense flux tubes, implying a systematic bias caused by our heuristic magnetisation criterion $\bar{\kappa}\bar{r}_{g}\sim 1$. Our methodology for modelling the conditional and unconditional data, as well as addressing this bias, is presented in Section 3.
@@ 3 Stochastic transport model
Motivated by our previous findings, we present a simplified stochastic model for the transport of charged test particles, resulting from a competition between magnetised and unmagnetised behaviour. We assume that magnetised behaviour is dictated by the most intense flux tubes which extend up to the correlation scale of the turbulence $\lambda_{\mathrm{fl}}\lesssim L_{u}$. Particles are strictly bound to those field lines, but may undergo pitch-angle scattering with rate $\theta_{\mu}$. This scattering represents large-scale mirror interactions on the scale of the flux tubes, and thus, its MFP $\lambda_{\mu}\sim c/\theta_{\mu}$ is expected to be long and comparable to the flux tube scale, i.e. $\lambda_{\mu}\sim\lambda_{\mathrm{fl}}$. Particles remain in the magnetised transport state for some mean duration $t_{\mu}^{\ast}$, which represents the mean time until a scattering event due to sharp field-line curvature occurs.
However, unmagnetised behaviour results from an interplay between the intrinsic particle momentum and random scattering due to highly tangled field lines, where no well-defined mean field is seen by the particle, so we expect a relatively short MFP and high scattering rate $\theta_{\mathrm{scatter}}\sim c/\lambda_{\mathrm{scatter}}$. Particles wander around the domain for some mean duration $t^{\ast}_{\mathrm{scatter}}$ until they encounter a strong flux tube and become magnetised again.
@@
The mean durations parametrise escape-time probability distributions $p_{\mu}(t|t^{\ast}_{\mu})$ and $p_{\mathrm{scatter}}(t|t^{\ast}_{\mathrm{scatter}})$ for both transport modes. They may be related to the distribution of field-line curvature, and from the dominance of low curvatures, we expect the ordering
@@
We assume here that each kind of motion (specifically field-line wandering, magnetised mirroring and unmagnetised scattering) is sufficiently characterised by its respective MFP $\lambda$ and thus adequately modelled by a Langevin equation, where both initial ballistic and asymptotic diffusive behaviour are resolved. The emergent combined transport behaviour can then be studied as a competition between the involved MFPs and mean durations. Further, the combination of two simple Langevin-like descriptions readily leads to compound subdiffusion and thus, a plausible explanation for the transient subdiffusion shown by the test particles. We emphasise that qualitative descriptions are linked to the relative sizes of the various MFPs, for instance, “coherent motion” or “large-scale mirroring” correspond to $\lambda_{\mu}\sim L_{u}$, while “chaotic motion” corresponds to $\lambda_{\mathrm{scatter}}\ll L_{u}$.